OpenStudy (taepoohwae):

find the area of the triangle with base 1 3/4 yards and height 4/5 yard.

OpenStudy (ajprincess):

Area of the triangle = 1/2*b*h b-base h-height Does that help @taepoohwae

OpenStudy (anonymous):

The Formula for area of triangle is : \[Area = \frac{1}{2} \times Base \times Height\]

OpenStudy (anonymous):

As you can see Base is in Mixed fraction so you can convert it into Normal Fraction.. \[a \frac{b}{c} = \frac{a \cdot c + b}{c}\]

OpenStudy (anonymous):

So, you are given with Height and Base, so you have to just plug in the values and then Find the Area of Triangle by just calculating..

OpenStudy (anonymous):

\[\frac{ 1 }{ 2}*base*height\]\[\frac{ 1 }{ 2 }*\frac{ 7 }{ 4 }*\frac{ 4 }{ 5 }\]\[\frac{ 7 }{ 10 }\]